# Data-driven algorithm design using neural networks with applications to branch-and-cut

### v1

@@ Appendix A Proofs of Main Results
using the notation from Section 2.2. We wish apply Lemma B.2 on the parameterized function class given by $h$ with $\mathcal{P}=\mathbb{R}^{W}$. Accordingly, we need to find a function $R:\mathbb{N}\to\mathbb{N}$ such that for any natural number $t>W$, and any $I_{1},\ldots,I_{t}$, the function $\mathbf{w}\mapsto(h(I_{1},\mathbf{w}),\ldots,h(I_{t},\mathbf{w}))$ is a piecewise polynomial function on $\mathbb{R}^{W}$ with at most $R(t)$ pieces.
@@
By Lemma B.4, there exist a decomposition of $\mathcal{W}^{\prime}$ into at most $\left(\frac{etU}{W^{\prime}}\right)^{W^{\prime}}$ regions, where $W^{\prime}=W-\ell w_{L}$ is the number of parameters determining the space $\mathcal{W}^{\prime}$, such that within each region, the output of the final hidden layer of the neural network is constant (as a function of the parameters in the region) for each input $\operatorname{Enc}(I_{j})$, $j\in[t]$.
@@
have the same signs, then $h(I_{j},(\mathbf{w},A^{L+1}))=S(I_{j},\varphi^{N^{\operatorname{sgn}},\operatorname{Sigmoid}}_{\mathbf{w},A^{L+1}}(I_{j}))$ is a polynomial of degree at most $\lambda.$ By the observations made above, the functions $\psi^{j}_{1}(T(\operatorname{Sigmoid}(A^{L+1}\mathbf{z}^{j}))),\ldots,\psi^{j}_{\Gamma}(T(\operatorname{Sigmoid}(A^{L+1}\mathbf{z}^{j})))$ are rational functions, i.e., ratios of polynomials, in the transformed parameters $\theta_{ki}$ and the numerators and denominators of these rational functions have degrees bounded by $2w_{L}\gamma$. Since $\operatorname{sgn}\left(\frac{P}{Q}\right)=\operatorname{sgn}(PQ)$ for any multivariate polynomials $P$, $Q$ (whenever the denominator is nonzero), we can bound the total number of sign patterns for $\psi^{j}_{1}(T(\operatorname{Sigmoid}(A^{L+1}\mathbf{z}^{j}))),\ldots,\psi^{j}_{\Gamma}(T(\operatorname{Sigmoid}(A^{L+1}\mathbf{z}^{j})))$ using Lemma B.1 where the polynomials defining the regions have degree at most $3w_{L}\gamma$ on the transformed parameters $\theta_{ki}$. We have to consider all the functions $\psi^{j}_{1}(T(\operatorname{Sigmoid}(A^{L+1}\mathbf{z}^{j}))),\ldots,\psi^{j}_{\Gamma}(T(\operatorname{Sigmoid}(A^{L+1}\mathbf{z}^{j})))$ for $j\in[t]$, giving us a total of $t\Gamma$ rational functions. Thus, an application of Lemma B.1 gives us a decomposition of ×$\mathbb{R}^{\ell\times w_{L}}$ into at most
@@
regions, such that within each region $h(I_{j},(\mathbf{w},A^{L+1}))$, as a function of $(\mathbf{w},A^{L+1})$, is a polynomial of degree at most $\lambda$, for every $j\in[t]$. Moreover, note that within each such region, $h(I_{j},(\mathbf{w},A^{L+1}))$ depends only on $A^{L+1}$. Applying Lemma B.2, $\operatorname{Pdim}\left(\mathcal{F}^{S}_{N^{\operatorname{sgn}},\operatorname{Sigmoid}}\right)$ is bounded by the largest $t\in\mathbb{N}$ such that
@@
As we only need to derive an upper bound for the pseudo-dimension, we can loosen the inequality above using the inequality $\log x\leq\frac{1}{c}x+\log(c/e)$ for real numbers $x,c>0$:
$\frac{1}{2}(t-2)$ $\leq W^{\prime}\left(\frac{etU/W^{\prime}}{8eU}+\log(8U)\right)+\ell w_{L}\left(\frac{6et\gamma\Gamma/\ell}{48e\gamma\Gamma w_{L}}+\log(48\gamma\Gamma w_{L})\right)+\ell w_{L}\left(\frac{2et(\lambda+1)/(\ell w_{L})}{16e(\lambda+1)}+\log(16(\lambda+1))\right)$
@@
using the notation from Section 2.2. We wish apply Lemma B.2 on the parameterized function class given by $h$ with $\mathcal{P}=\mathbb{R}^{W}$. Accordingly, we need to find a function $R:\mathbb{N}\to\mathbb{N}$ such that for any natural number $t>LW$, and any $I_{1},\ldots,I_{t}$, the function $\mathbf{w}\mapsto(h(I_{1},\mathbf{w}),\ldots,h(I_{t},\mathbf{w}))$ is a piecewise polynomial function on $\mathbb{R}^{W}$ with at most $R(t)$ pieces.
Note that $\varphi^{\operatorname{ReLU},\operatorname{CReLU}}_{\mathbf{w}}(I)$ can be seen as the output of a neural network with ReLU activations and architecture $[d,w_{1},\ldots,w_{L},2\ell,\ell]$, where the final linear function is the fixed function $T(\mathbf{u})_{k}=(\tau_{k}-\eta_{k})\mathbf{u}_{k}$. This is because $\operatorname{CReLU}(x)=\operatorname{ReLU}(x)-\operatorname{ReLU}(x-1)$ can be simulated using two $\operatorname{ReLU}$ neurons. Applying Lemma B.3, this implies that given $t\in\mathbb{N}$ and $I_{1},\ldots,I_{t}\in\mathcal{I}$, there are
@@
have the same signs, then $h(I_{j},\mathbf{w})=S(I_{j},\varphi^{\operatorname{ReLU},\operatorname{CReLU}}_{\mathbf{w}}(I_{j}))$ is a polynomial of degree at most $\lambda(L+2).$ The functions $\psi^{j}_{1}(\varphi^{\operatorname{ReLU},\operatorname{CReLU}}_{\mathbf{w}}(I_{j})),\ldots,\psi^{j}_{\Gamma}(\varphi^{\operatorname{ReLU},\operatorname{CReLU}}_{\mathbf{w}}(I_{j}))$ are polynomials of degree at most $\gamma(L+2)$. Considering all these polynomials for $j\in[t]$, by Lemma B.1, each $\mathcal{W}_{i}\subseteq\mathbb{R}^{W}$ from the decomposition above can be further decomposed into at most $2\left(\frac{2et\Gamma\gamma(L+2)}{W}\right)^{W}$ regions such that for all $\mathbf{w}$ in such a region, $h(I_{j},\mathbf{w})$ is a polynomial function of $\mathbf{w}$ of degree at most $\lambda(L+2)$.
@@
regions such that within each region, $h(I_{j},\mathbf{w})$ is a polynomial function of $\mathbf{w}$ of degree at most $\lambda(L+2)$. Applying Lemma B.2, $\operatorname{Pdim}\left(\mathcal{F}^{S}_{N^{\operatorname{ReLU}},\operatorname{CReLU}}\right)$ is bounded by the largest $t\in\mathbb{N}$ such that
@@
$\frac{1}{2}(t-1)$ $\leq L+2+(L+1)W\log\left(\frac{2et(U+2\ell)}{W}\right)+W\log\left(\frac{2et\Gamma\gamma(L+2)}{W}\right)+W\log\left(\frac{2et(\lambda+1)(L+2)}{W}\right)$
$\leq L+2+\frac{1}{8}t+(L+1)W\log(16(U+2\ell))+\frac{1}{8}t+W\log(16\Gamma\gamma(L+2))+\frac{1}{8}t+W\log(16(\lambda+1)(L+2))$
@@
then it’s not hard to see that
@@
We introduce an auxiliary function $f:\mathbb{R}^{r}\rightarrow\{\mathbf{p}_{1},\dots,\mathbf{p}_{r}\}$ given by $f(\mathbf{x})=\mathbf{p}_{\operatorname{arg\,max}_{i\in[r]}\mathbf{x}_{i}}$, and let $S^{\prime}:\mathcal{I}\times\mathbb{R}^{r}\rightarrow\mathbb{R}$ as
@@
Notice that within each region decomposed by at most $\binom{r}{2}$ hyperplanes $\{\mathbf{x}\in\mathbb{R}^{r}:\mathbf{x}_{i}=\mathbf{x}_{j}\}$, with $(i,j)\in[r]\times[r]$ and $i\neq j$, the largest coordinate of $\mathbf{x}$ is unchanged. Therefore, for any fixed $I\in\mathcal{I}$, the new score function $S^{\prime}(I,\cdot)$ remains constant in each of these regions. Then a direct application of Theorem 2.5 and Theorem 2.6 to $S^{\prime}$ yields the desired result. ∎
Appendix B Auxiliary Lemmas
Lemma B.1.
[Theorem 5.5 in Matousek [1999], Lemma 17 in Bartlett et al. [2019], Lemma 3.3 in Anthony et al. [1999]] Let $\mathcal{P}\subseteq\mathbb{R}^{\ell}$ and let $f_{1},\ldots,f_{t}:\mathbb{R}^{\ell}\to\mathbb{R}$ with $t\geq\ell$ be functions that are polynomials of degree $m$ when restricted to $\mathcal{P}$. Then
|$|\{\operatorname{sgn}(f_{1}(\mathbf{p})),\ldots,\operatorname{sgn}(f_{t}(\mathbf{p})):\mathbf{p}\in\mathcal{P}\}|$ $\leq 2\left(\frac{2etm}{\ell}\right)^{\ell},$ $m\geq 2,$
|$|\{\operatorname{sgn}(f_{1}(\mathbf{p})),\ldots,\operatorname{sgn}(f_{t}(\mathbf{p})):\mathbf{p}\in\mathcal{P}\}|$ $=1,$ $m=0.$
Lemma B.2.
$\mathbf{p}\mapsto(h(I_{1},\mathbf{p}),\ldots,h(I_{t},\mathbf{p}))$
Proof.
Within each $\mathcal{P}_{i}$, $h(I_{j},\mathbf{p})-s_{j}$ is a polynomial in $p$ for every $j=1,\ldots,t$, given the hypothesis. Applying Lemma B.1,
Summing over the different $\mathcal{P}_{i}$, $i\in[R(t)]$, we obtain that
Lemma B.3.
Proof.
Follows from Section 4 in Bartlett et al. [2019]. ∎
Lemma B.4.
Proof.
Fix a $t\in\mathbb{N}$ and $\mathbf{x}^{1},\ldots,\mathbf{x}^{t}\in\mathbb{R}^{w_{0}}$. Consider a neuron in the first hidden layer. The output of this neuron on the input $\mathbf{x}^{j}$ is $\operatorname{sgn}(\langle\mathbf{a},\mathbf{x}^{j}\rangle+b)$, where $\mathbf{a}\in\mathbb{R}^{w_{0}},b\in\mathbb{R}$ are the weights associated with this neuron. As a function of these weights, this is a linear function, i.e., a polynomial function of degree 1. Applying Lemma B.1, there are at most $\left(\frac{et}{w_{0}+1}\right)^{w_{0}+1}$ regions in the space $(\mathbf{a},b)\in\mathbb{R}^{w_{0}}\times\mathbb{R}$ such that within each region, the output of this neuron is constant. Applying the reasoning for the $w_{1}$ neurons in the first layer, we obtain that the output of the first hidden layer (as a $0/1$ vector in $\mathbb{R}^{w_{1}}$) is piecewise constant as a function of the parameters of the first layer, with at most $\left(\frac{et}{w_{0}+1}\right)^{(w_{0}+1)w_{1}}$ pieces. For a fixed output of the first hidden layer (which is a vector in $\mathbb{R}^{w_{1}}$), we can apply the same reasoning and partition space of weights of the second layer into $\left(\frac{et}{w_{1}+1}\right)^{(w_{1}+1)w_{2}}$ regions where the output of the second hidden layer is constant. Applying this argument iteratively across the hidden layers, and using Young’s inequality, we deduce that a decomposition exists for $\mathbb{R}^{W}$ with at most
$=$ $\left(\frac{etU}{W^{\prime}}\right)^{W^{\prime}}$

### v2

@@ Appendix B Proofs of main results
using the notation from Section 2.2. We wish apply Lemma A.3 on the parameterized function class given by $h$ with $\mathcal{P}=\mathbb{R}^{W}$. Accordingly, we need to find a function $R:\mathbb{N}\to\mathbb{N}$ such that for any natural number $t>W$, and any $I_{1},\ldots,I_{t}$, the function $\mathbf{w}\mapsto(h(I_{1},\mathbf{w}),\ldots,h(I_{t},\mathbf{w}))$ is a piecewise polynomial function on $\mathbb{R}^{W}$ with at most $R(t)$ pieces.
@@
By Lemma A.5, there exist a decomposition of $\mathcal{W}^{\prime}$ into at most $\left(\frac{etU}{W^{\prime}}\right)^{W^{\prime}}$ regions, where $W^{\prime}=W-\ell w_{L}$ is the number of parameters determining the space $\mathcal{W}^{\prime}$, such that within each region, the output of the final hidden layer of the neural network is constant (as a function of the parameters in the region) for each input $\operatorname{Enc}(I_{j})$, $j\in[t]$.
@@
have the same signs, then $h(I_{j},(\mathbf{w},A^{L+1}))=S(I_{j},\varphi^{N^{\operatorname{sgn}},\operatorname{Sigmoid}}_{\mathbf{w},A^{L+1}}(I_{j}))$ is a polynomial of degree at most $\lambda.$ By the observations made above, the functions $\psi^{j}_{1}(T(\operatorname{Sigmoid}(A^{L+1}\mathbf{z}^{j}))),\ldots,\psi^{j}_{\Gamma}(T(\operatorname{Sigmoid}(A^{L+1}\mathbf{z}^{j})))$ are rational functions, i.e., ratios of polynomials, in the transformed parameters $\theta_{ki}$ and the numerators and denominators of these rational functions have degrees bounded by $2w_{L}\gamma$. Since $\operatorname{sgn}\left(\frac{P}{Q}\right)=\operatorname{sgn}(PQ)$ for any multivariate polynomials $P$, $Q$ (whenever the denominator is nonzero), we can bound the total number of sign patterns for $\psi^{j}_{1}(T(\operatorname{Sigmoid}(A^{L+1}\mathbf{z}^{j}))),\ldots,\psi^{j}_{\Gamma}(T(\operatorname{Sigmoid}(A^{L+1}\mathbf{z}^{j})))$ using Lemma A.2 where the polynomials defining the regions have degree at most $3w_{L}\gamma$ on the transformed parameters $\theta_{ki}$. We have to consider all the functions $\psi^{j}_{1}(T(\operatorname{Sigmoid}(A^{L+1}\mathbf{z}^{j}))),\ldots,\psi^{j}_{\Gamma}(T(\operatorname{Sigmoid}(A^{L+1}\mathbf{z}^{j})))$ for $j\in[t]$, giving us a total of $t\Gamma$ rational functions. Thus, an application of Lemma A.2 gives us a decomposition of ×$\mathbb{R}^{\ell\times w_{L}}$ into at most
@@
regions, such that within each region $h(I_{j},(\mathbf{w},A^{L+1}))$, as a function of $(\mathbf{w},A^{L+1})$, is a polynomial of degree at most $\lambda$, for every $j\in[t]$. Moreover, note that within each such region, $h(I_{j},(\mathbf{w},A^{L+1}))$ depends only on $A^{L+1}$. Applying Lemma A.3, $\operatorname{Pdim}\left(\mathcal{F}^{S}_{N^{\operatorname{sgn}},\operatorname{Sigmoid}}\right)$ is bounded by the largest $t\in\mathbb{N}$ such that
@@
As we only need to derive an upper bound for the pseudo-dimension, we can loosen the inequality above using the inequality (3) in Lemma A.1:
$\frac{1}{2}(t-2)\leq$ $W^{\prime}\left(\frac{etU/W^{\prime}}{8eU}+\log(8U)\right)+\ell w_{L}\left(\frac{6et\gamma\Gamma/\ell}{48e\gamma\Gamma w_{L}}+\log(48\gamma\Gamma w_{L})\right)$
$+\ell w_{L}\left(\frac{2et(\lambda+1)/(\ell w_{L})}{16e(\lambda+1)}+\log(16(\lambda+1))\right)$
@@
using the notation from Section 2.2. We wish apply Lemma A.3 on the parameterized function class given by $h$ with $\mathcal{P}=\mathbb{R}^{W}$. Accordingly, we need to find a function $R:\mathbb{N}\to\mathbb{N}$ such that for any natural number $t>LW$, and any $I_{1},\ldots,I_{t}$, the function $\mathbf{w}\mapsto(h(I_{1},\mathbf{w}),\ldots,h(I_{t},\mathbf{w}))$ is a piecewise polynomial function on $\mathbb{R}^{W}$ with at most $R(t)$ pieces.
Note that $\varphi^{\operatorname{ReLU},\operatorname{CReLU}}_{\mathbf{w}}(I)$ can be seen as the output of a neural network with ReLU activations and architecture $[d,w_{1},\ldots,w_{L},2\ell,\ell]$, where the final linear function is the fixed function $T(\mathbf{u})_{k}=(\tau_{k}-\eta_{k})\mathbf{u}_{k}$. This is because $\operatorname{CReLU}(x)=\operatorname{ReLU}(x)-\operatorname{ReLU}(x-1)$ can be simulated using two $\operatorname{ReLU}$ neurons. Applying Lemma A.4, this implies that given $t\in\mathbb{N}$ and $I_{1},\ldots,I_{t}\in\mathcal{I}$, there are
@@
have the same signs, then $h(I_{j},\mathbf{w})=S(I_{j},\varphi^{\operatorname{ReLU},\operatorname{CReLU}}_{\mathbf{w}}(I_{j}))$ is a polynomial of degree at most $\lambda(L+2).$ The functions $\psi^{j}_{1}(\varphi^{\operatorname{ReLU},\operatorname{CReLU}}_{\mathbf{w}}(I_{j})),\ldots,\psi^{j}_{\Gamma}(\varphi^{\operatorname{ReLU},\operatorname{CReLU}}_{\mathbf{w}}(I_{j}))$ are polynomials of degree at most $\gamma(L+2)$. Considering all these polynomials for $j\in[t]$, by Lemma A.2, each $\mathcal{W}_{i}\subseteq\mathbb{R}^{W}$ from the decomposition above can be further decomposed into at most $2\left(\frac{2et\Gamma\gamma(L+2)}{W}\right)^{W}$ regions such that for all $\mathbf{w}$ in such a region, $h(I_{j},\mathbf{w})$ is a polynomial function of $\mathbf{w}$ of degree at most $\lambda(L+2)$.
@@
regions such that within each region, $h(I_{j},\mathbf{w})$ is a polynomial function of $\mathbf{w}$ of degree at most $\lambda(L+2)$. Applying Lemma A.3, $\operatorname{Pdim}\left(\mathcal{F}^{S}_{N^{\operatorname{ReLU}},\operatorname{CReLU}}\right)$ is bounded by the largest $t\in\mathbb{N}$ such that
@@
$\frac{1}{2}(t-1)\leq$ $L+2+(L+1)W\log\left(\frac{2et(U+2\ell)}{W}\right)+W\log\left(\frac{2et\Gamma\gamma(L+2)}{W}\right)$
$+W\log\left(\frac{2et(\lambda+1)(L+2)}{W}\right)$
$\leq$ $L+2+\frac{1}{8}t+(L+1)W\log(16(U+2\ell))+\frac{1}{8}t+W\log(16\Gamma\gamma(L+2))$
$+\frac{1}{8}t+W\log(16(\lambda+1)(L+2))$
@@
where the inequality (3) in Lemma A.1 is applied in the second line. Then it’s not hard to see that
@@
We introduce an auxiliary function $f:\mathbb{R}^{r}\rightarrow\{\mathbf{p}_{1},\dots,\mathbf{p}_{r}\}$ given by $f(\mathbf{x})=\mathbf{p}_{\operatorname{arg\,max}_{i\in[r]}\mathbf{x}_{i}}$, and let $S^{\prime}:\mathcal{I}\times\mathbb{R}^{r}\rightarrow\mathbb{R}$ be
@@
There exists a decomposition of the $\mathbb{R}^{r}$ space obtained by at most $\frac{r(r-1)}{2}$ hyperplanes
$\left\{\mathbf{x}\in\mathbb{R}^{r}:\mathbf{x}_{i}=\mathbf{x}_{j}\right\},\quad\forall(i,j)\in[r]\times[r],i\neq j.$
Within each decomposed region, the largest coordinate of $\mathbf{x}$ is unchanged. Therefore, for any fixed $I\in\mathcal{I}$, the new score function $S^{\prime}(I,\cdot)$ remains constant in each of these regions. Then a direct application of Theorem 2.5 and Theorem 2.6 to $S^{\prime}$ yields the desired result. ∎
Proof of Proposition 3.7.
The proof of Proposition 3.7 is analogous to the proof of Theorem 4.1 in Balcan et al. [2021d]. For any fixed instance $I\in\mathcal{I}$, let the set of candidate cutting planes be $\mathcal{C}=\{c_{1},\dots,c_{|\mathcal{C}|}\}$. Comparing the overall scores for all cuts introduces the following hyperplanes:
$\sum_{i=1}^{\ell}\bm{\mu}_{i}\operatorname{Score}_{i}(c_{j},I)=\sum_{i=1}^{\ell}\bm{\mu}_{i}\operatorname{Score}_{i}(c_{k},I),\quad\forall j,k\in[|\mathcal{C}|],j\neq k.$
There are at most $\frac{|\mathcal{C}|(|\mathcal{C}|-1)}{2}$ hyperplanes decomposing the $\bm{\mu}$ space $[0,1]^{\ell}$. Within each region defined by these hyperplanes, the selected cut remains the same, so the branch-and-cut tree size is constant. This proves that $f_{S}(I,\bm{\mu})$ is a piecewise constant function on $\bm{\mu}$, for any fixed $I$. We then apply Theorem 2.5 and Theorem 2.6 to derive the pseudo-dimension bounds. ∎